A truck moving at a ratr of 50 m/s sounds its horn with a frequency of 800 Hz. Find the frequency heard by the stationary observer, as the truck approaches him. The ambient temperature is about 40°C.


1
Expert's answer
2021-02-17T11:08:42-0500

Find the speed of sound at 40°C:


"c=331\sqrt{1+\frac{40}{273}}=354\text{ m/s}."

The frequency heard by the stationary observer, as the truck approaches:


"f_r=\frac{f_s}{1-v/c}=932\text{ Hz}."